ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 176, 2010

Alesker, S. and M. Verbitsky Quaternionic Monge–Ampère equation and Calabi problem for HKT-manifolds

109

Bourgain, J. Estimates on polynomial exponential sums  221

Isabel Cortez, M. I. and J. Rivera-Letelier Invariant measures of minimal post-critical sets of logistic maps

157

del Río, Á. See A. Herman Di Vincenzo, O. M. and V. Nardozza On a Regev–Seeman con-

jecture about ℤ_{2}-graded tensor products

43

Gorodnik, A., H. Oh and N. Shah Strong wavefront lemma and counting lattice points in sectors

419

Han, Zhigang The bounded isometry conjecture for the Kodaira-Thurston manifold and the 4-torus

285

Harbater, D. and J. Hartmann Patching over fields  61

Hartmann, J. See D. Harbater Herman, A., G. Olteanu and Á. del Río The gap between the

Schur group and the subgroup generated by cyclic cyclotomic algebras

401

Hou, J. and L. Huang Maps completely preserving idempotents and maps completely preserving square-zero operators

363

Huang, L. See J. Hou Konyagin, S. V. and V. F. Lev On the number of popular differ-

ences

269

Krieger, W. and B. Weiss On g-measures in symbolic dynamics  1

Lee, C. W. and M. M. Menzel A generalized sewing construction for polytopes

241

Lev, V. F. See S. V. Konyagin

Lotz, H. P. Weak∗ convergence in the dual of Weak L^{p}  209

Lusky, W. and J. Taskinen On weighted spaces of holomorphic functions of several variables

381

Menzel, M. M. See C. W. Lee Nachmias, A. and Y. Peres The critical random graph, with mar-

tingales

29

Nardozza, V. See O. M. Di Vincenzo Oh, H. See A. Gorodnik Olteanu, G. See A. Herman Peres, Y. See A. Nachmias Rivera-Letelier, J. See M. I. Cortez Schraen, B. Représentations p-adiques de GL_{2}(L) et catégories

dérivées

307

Shah, N. See A. Gorodnik Steinberg, B. Maximal subgroups of the minimal ideal of a free profi-

nite monoid are free

139

Taskinen, J. See W. Lusky Veliev, O. A. On the Nonself-adjoint Ordinary Differential Operators

with Periodic Boundary Conditions

195

Verbitsky, M. See S. Alesker Weiss, B. See W. Krieger